Find the derivative. Simplify if possible.
f(t) = (sech(9e power t))2
1
Expert's answer
2011-02-28T08:31:57-0500
sech(z) = 2 (ez + e-z)-1;[sech(z)]' = - 2 (ez - e-z)/(ez + e-z)2 = - sech(z) tanh(z);
f'(t) = 2 [sech (9et)]' sech(9et) = - 2 [9et]' sech2 (9et) tanh(9et) = - 18 et sech2 (9et) tanh(9et)